The initial concentration of cell culture is 500 cells per mL. After 3 hours, the of cells in one mL is 2,56,000. calculate the generation time of cells
1
Expert's answer
2020-06-22T13:29:49-0400

The generation time is the time interval required for the cells (or population) to divide.

G (generation time) =t (time, in minutes)/n(number of generations)

G = t/n

t = time interval in minutes = 180 minutes

B = number of bacteria at the beginning of a time interval = 500 cells 

b = number of bacteria at the end of the time interval = 2,56,000 cells

n = number of generations (number of times the cell population doubles during the time interval)

b = B x 2n (This equation is an expression of growth by binary fission)

Solve for n:

logb = logB + nlog2

n = logb - logB

          log2

n = logb - logB

          .301

n = 3.3 logb/B

G = t/n

Solve for G

G =       t

      3.3 log b/B

G =       180

      3.3 log 256000 –log500

=180 / 8.943

G = 20.12 minutes per generation